1

a) List four (4) segments of the circular flow.         

b) Name and explain the importance of two markets in a circular flow

2 Suppose the following demand and supply functions are given as:

 Qs = 4 + 4P and Qd =40 - 8p.

a)    Calculate the market clearing price and quantity for the above market and show the results on the graph.                                                                                            [05]

b)    Suppose the price increase by 100%, what will happen to the quantity demanded and quantity supplied. Show the changes on a new graph.                                     [05]

3 Name and explain two (2) types of disequilibrium in the market.

4 Differentiate between movement along and a shift in demand curve.                        [02]

5 List any two (2) determinants of demand and explain how each can change the demand curve.  

                                                      


1
Expert's answer
2022-02-28T11:41:38-0500

1

a) Individuals, firms, market for goods and services, market for factors of production.

b)The two markets in the circular flow help to maintain a the flow of goods and services and money around the economy. This is mostly done through the firms and households. Money flows form the firms to workers (members of households) in form of wages and salaries while mo ey flows back to them in form of payment for purchasing those goods that have been produced by them.


2

a)"E=Q_d=Q_s"

Where Qd = 40-8P and Qs = 4+4P

"40-8P=4+4P"

"P^*=3"

"Q^*=4+4(3)=16"


b) if price increases by "100\%" , "{P^{*}{_n}}=6"

Then quantity demanded and quantity supplied will increase decrease by 150%. "{Q_n}^{*}=-8". In essence, nobody buys anymore because the prices are way too high.


3. Surpluses and shortages.

Surpluses occurs when the price is above the equilibrium price in the market. This is a case of excess supply. Suppliers wants to produce and sell more.

Shortages, or excess demand, occur when the price is below the equilibrium price in the market. This leads consumers to want to buy more at that price


4) A movement along the demand curve is caused by changes in the price of a commodity while other factors affecting demand are constant. On the other hand, shift in demand is caused by the changes in other factors affecting demand except the commodity's own price. This may be an outward or inward shift.


5) i) Price

ii) income


The price of a particular commodity can affect its demand curve. An increase in the price of such commodity can lead to consumers buying less of such goods(quantity demanded decreases) while a decrease in the price of such commodity can lead consumers to buy more of it thereby leading to an increase in quantity demanded.


Income plays a major role in the demand for any goods. An income increase will lead to an increase in the demand for any commodity given that such commodity is a normal good and vice versa. If such commodity where an inferior good, as income increases, the the consumption of the good decreases leading to a lower or lesser demand for the good.